Problem 2. The firm expects that its production function in the next period will be given by Fe (KL) 10VKL. The firm employs 25 workers. The real interest rate equals 5% and the depreciation rate equals 20%. The price for the firms products is 1 EUR while the price of capital goods is 1O EUR Find the user cost of capital. Find net investments, assuming that initially the firm has a capital stock of 88. Modify your answer to part (a), assuming there is a 10% tax on the firms profits. a) b) The firm considers hiring additional workers. How would this decision affect your answer to part (a)?
Answer #1

A)

Real(user) cost of capital = (Pk/P) (r + δ )

Pk = price of capital = 10

P = general price level = 1

r = real interest rate = 5% = 0.05

δ = depreciation raten = 20% = 0.20

Hence, R, = (10/1)(0.05 0.20) 2.5

At this user cost of capital, ideal level of capital will be;

MPK (Marginal Product of Capital) = rental cost of capital

MPK-5K-12 L12 = 2.5

= MPK = 5K-1122512 = 2.5

= К 100

Net Investment = Change in capital * price of capital good

Net Investment = (100 - 88 )* 10= 120 EUR.

B)The rental cost of capital is affected by taxes. A tax on corporate profit would discourage business investment. Cost of capital tends to rise with the tax rate.

As we can see in the equation of MPK; The greater the amount of labor employed the higher is the rental cost of capital.